April 2004

Sunday Monday Tuesday Wednesday Thursday Friday Saturday
 
 
 
 
 
 
1
LAN fiber termination completed.

Begin staging bare steel plates, moving them out of New Muon and stacking them on the floor of MSB.

Electricians continue in the Hall and in the Absorber area.

Pipefitters continue in Ab area.
2
Continue staging bare steel plates to MSB.

Pipefitters finish all work within the Ab Enclosure.  Move to the skid area and continue work there.

3

 
 
 
 
4
Projected Work:
• Begin installing racks in Hall, with associated cabling
• Continue electrical work in Hall & Ab areas
• Continue pipefitters in Absorber Area
5
Computing Div begins LAN hardware installation.

Plane 281 checked for light leaks. OK.

Plane 276 (instrumented) moved from NM to MSB. This plane was already disconnected and on a strongback, so it could be reversed on the NM stand and light-leak checked (leak found and fixed).

First 4 racks moved from NM to MSB, all FE type racks (with phototubes). Moving procedure and vibration absorbers appear to work as expected.

Misc. electrical work in Hall & Ab area continues.

Pipefitters continue in Ab Skid area.
6
LAN hardware installation complete. Hall LAN is operational. Also have a couple of DHCP ports in MSB.

Crew training and practice for plane installation. Includes fill-in techs in addition to the regular installation crew. Install planes 280 & 279 (bare steel).

Vulcan system operational. Survey crew is also present to cross check locations of steel and coil collar.

First 4 racks, moved to MSB yesterday, go underground.

Misc. electrical work in Hall & Ab area continues.

Pipefitters continue in Ab Skid area.
7
Crew training and practice for plane installation continues. Install planes 278 & 277 (bare steel), & 276 (instrumented).

Misc. electrical work in Hall & Ab area continues.

Pipefitters continue in Ab Skid area.
8
Readout Test of 9 planes at New Muon decommissioned. Racks prepped for moving underground.

Stage  planes to MSB, including the next instrumented one, 271.

Prepare more bare steel plates at New Muon.

Move some racks underground.

Misc. electrical work in Hall & Ab area continues.

Pipefitters continue in Ab Skid area.
9
Move more racks underground. A total of 24 now underground - all are staged on the Hall floor, and not in final locations.

Move Plane commissioning equipment (cable connection hardware & tools) underground.

Coil Power Supply and reversing switch moved underground.

Cable and light-leak check Planes 281 & 276.

Misc. electrical work in Hall & Ab area continues.

Pipefitters continue in Ab Skid area.
10
Techs move 24 racks into place alongside the detector.
11
Projected Work:
• Begin 5 Plane per day Installation
• Continue installing racks in Hall
• Complete electrical work in Detector Hall
• Test pumpdown of DK vacuum - during the weekend (NOT) 4/17-18
• Continue pipefitters in Skid Area
12
Five planes installed - 275, 274, 273, 272, 271. Plane 271 has cables attached and light-leak checked by 4pm.

Electricians work on power for Muon Detector readout racks.

Pipefitters continue in Ab area - work there is nearly complete.
13
Five planes installed - 270, 269, 268, 267, 266. Plane 266 has cables attached and light-leak checked by 5pm.

Problem with tightening fixture on plane transport cart - it is tearing itself apart.

Electricians complete work on power for Muon Detector readout racks.

Pipefitters move to MINOS Hall to work on Hall LCW.
14
No planes installed. Replace plane transport cart tightening fixture with a better design. Also start to remove guardrail on upper rack level, as it interferes with the fiber cable tray placement.

No shaft access. Sump pump #2 re-installed, which involves removing sump covers at shaft base.

Electricians work on AC cord extensions on all FE racks.

Painters start on MSB floor "stay clear" areas in front of power panels.

Pipefitters continue with MINOS LCW system in Hall.
15
Five planes installed - 265, 264, 263, 262, 261. Plane 261 has cables attached and light-leak checked by 6pm.

Elevator door breaks at 5pm.

Electricians continue with FE rack AC cords.

Pipefitters complete work.  Tools and materials rigged out and sent to Target Hall.
16
No planes installed.  Elevator repair. Limited to 6 persons underground.

DAQ racks are ready to move underground, so these plus the Clock rack are delivered and set into place.

Continue with upper level guardrail removal and installation of replacements which do not interfere with fiber cable trays.

Electricians continue with FE rack AC cords.

Hadron Monitor support structure delivered, and stored in Muon Alcove #3.
17

 
 
 
 
18
Projected Work:
• Continue Plane Installation
• Move DAQ racks underground and begin installation of DAQ
• Complete readout rack installation
• No Absorber Inst work - making plans for installation sans forktruck
19
Sump #2 fails again.

Five planes installed - 260, 259, 258, 257, 256. Plane 256 has cables attached and light-leak checked by 6pm.

Commissioning Shifters work on readout cabling in addition to fiber readout cabling, all this week.

Electricians continue with FE rack AC cords.

Additional phones installed in Hall.
20
Five planes installed - 255, 254, 253, 252, 251. Plane 251 has cables attached and light-leak checked by 6pm.

Sump #2 pulled out at 3:30pm.  Found a broken wire, squashed during prior replacement. Shipped to manufacturer for checkout.

Electricians continue with FE rack AC cords.   Also install additional outlet on East Wall for a Controls Box.
21
Five planes installed - 250, 249, 248, 247, 246. Plane 246 has cables attached and light-leak checked by 6pm.

Electricians continue with FE rack AC cords.
22
Five planes installed - 245, 244, 243, 242, 241. Plane 241 has cables attached and light-leak checked by 6pm.

Electricians work on Absorber area Controls racks.

Begin floor painting of stay-clear areas in MINOS Hall.

Beams Controls installation work on Absorber area Controls racks.

Floor painting around Absorber area power panels.
23
Five planes installed - 240, 239, 238, 237, 236. Plane 236 has cables attached and light-leak checked by 6pm.

Continue floor painting in Absorber area and in MINOS Hall.

Beams Controls installation work in MINOS Hall Controls racks.
24

 
 
 
 
25
Projected Work:
• Continue Plane Installation
• Test pumpdown of DK vacuum - during the weekend  5/1-2
26
No planes installed.

Deliver and set into place 8 more racks (all FE V).

Deliver coil collars and catwalk supports to MINOS Hall.

Sufficient planes are installed to install the first catwalk section on top of the detector.
27
Five planes installed - 235, 234, 233, 232, 231. Plane 231 has cables attached and light-leak checked by 6pm.
28
Five planes installed - 230, 229, 228, 227, 226. Plane 226 has cables attached and light-leak checked by 6pm.

Last 4 detector readout racks delivered and set into place.  All Near Detector racks are now underground.

Readout cabling and DAQ commissioning work continues.
29
Five planes installed - 225, 224, 223, 222, 221. Plane 221 has cables attached and light-leak checked by 5pm.

Installation crew now starts at 6am. 5th plane is up by 12:30.  Commissioning shift begins at 2:30.

Readout cabling and DAQ commissioning work continues.

Electricians complete AC cord extension on BiRa AC trip boxes.
30
Five planes installed - 220, 219, 218, 217, 216. Plane 216 has cables attached and light-leak checked by 3:30pm.

Pump-down started on DK pipe at noon.  Reaches 1/2 atm by 4pm, at which point personnel are not allowed underground.

Rigging equipment for Absorber installation starts to arrive.